OpenStudy (anonymous):

Lucy borrowed 3$,500 at an apr of 10% for furniturre she will make 12 equal monthly payments. What will her monthly payments be?

OpenStudy (anonymous):

Amount = principal (1 + no. of year x rate) i.e A = P (1 + nr) A = 3500 (1 + 1 x 10%) A = 3500 (1 + 0.1) A = 3500 (1.1) A = 3850 for further help log on to www.teachon.tk

OpenStudy (anonymous):

Monthly pymt = $307.71